OpenStudy (anonymous):

identify the missing factor in (5a^4)(?)=5^3 a^8

OpenStudy (anonymous):

5^2a^4

OpenStudy (anonymous):

?=(5^3 a^8)/(5a^4)=5^2a^4

OpenStudy (anonymous):

thats not one of the multiple choice options

OpenStudy (anonymous):

so im guessing thats not the right answer

OpenStudy (anonymous):

25a^4 is not the ans????????

OpenStudy (anonymous):

oh yah actually it is.....can you explain 2 me again how to do it because i still dont quite unerstand how to get that answer

OpenStudy (anonymous):

5^2=25....my first ans was 5^2*a^4.....now i replaced 5^2 by 25

OpenStudy (anonymous):

wait so where are you getting 5^2

OpenStudy (anonymous):

?=(5^3 a^8)/(5a^4)=5^2a^4 here 5^3/5 gives me 5^2...